# CMSO-transducing tree-like graph decompositions

### v2

@@
Abstract
We show that given a graph $G$ we can $\operatorname{C}{\!\!}\operatorname{MSO}$-transduce its modular decomposition, its split decomposition and its bi-join decomposition. This improves results by Courcelle [Logical Methods in Computer Science, 2006] who gave such transductions using order-invariant $\operatorname{MSO}$, a strictly more expressive logic than $\operatorname{C}{\!\!}\operatorname{MSO}$. Our methods more generally yield $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transductions of the canonical decomposition of weakly-partitive set systems and weakly-bipartitive systems of bipartitions.
keywords:
@@ Abstract
category:
[…]
1 Introduction
@@
For a logic $\mathsf{L}$, an $\mathsf{L}$-transduction is a non-deterministic map from a class of relational structures to a new class of relational structures using $\mathsf{L}$-formulas. Transducing a tree-like decomposition is of particular interest. Notably, transducing a decomposition of a graph implies that any property that is definable using a decomposition, is also definable on graphs that admit such a decomposition. Moreover, tree-like decompositions can be often represented by labelled trees, for which the equivalence of recognisability by a tree automaton and definability in CMSO is well known [8]. Hence, it is an interesting question to consider what kind of graph decompositions can be transduced using $\mathsf{L}$-transductions for some extension $\mathsf{L}$ of $\operatorname{MSO}$.
@@
In a series of papers [8, 9, 10, 12], Courcelle investigated the relationship between the graph properties that can be defined in an extension of $\operatorname{MSO}$ and the graph properties that can be recognized by a tree automaton. In particular, for graphs of bounded treewidth, Courcelle’s theorem states that any property that is definable in the logic $\operatorname{MSO}$ with modulo counting predicate, denoted $\operatorname{C}{\!\!}\operatorname{MSO}$, can be recognized by a tree automaton [8]. Combining this result with the linear time algorithm for computing tree-decompositions [2], yields that $\operatorname{C}{\!\!}\operatorname{MSO}$ model-checking can be done in linear time on graphs of bounded treewidth. The converse statement – whether recognisability by a tree automaton implies definability in $\operatorname{C}{\!\!}\operatorname{MSO}$ on graphs of bounded treewidth – was conjectured by Courcelle in [8] and finally settled by Bojańczyk and Pilipczuk [5]. The key step to obtain this result is obtaining a tree-decomposition of a graph via an $\operatorname{MSO}$-transduction, a strategy which was initially proposed in [9] and is now standard.
@@
The obvious next question is whether an analogous result can be proved for graphs of bounded clique-width and for more general combinatorial objects, most notably, matroids representable over a fixed finite field and of bounded branch-width. Due to the above-mentioned strategy, the key challenge is to produce corresponding tree-like decompositions by $\operatorname{MSO}$-transduction. It is known that clique-width decompositions can be $\operatorname{MSO}$-transduced for graphs of bounded linear clique-width [4]. However, it is unknown if clique-width decompositions can be $\operatorname{MSO}$-transduced in general. In fact, this question remains open even for distance-hereditary graphs, which are precisely graphs of rank-width 1 (thus, of constant clique-width).
@@
Besides tree-decompositions, the problem of transducing cotrees, and in general hierarchical decompositions such as modular decompositions and split decompositions were considered in the literature [10, 11, 12, 14]. In [10], Courcelle provides transductions using order-invariant $\operatorname{MSO}$ for cographs and modular decompositions of graphs of bounded modular width. Order-invariant $\operatorname{MSO}$ allows the use of a linear order of the set of vertices and is more expressive than $\operatorname{C}{\!\!}\operatorname{MSO}$ [21]. The applicability of these transductions was later generalized using the framework of weakly-partitive set systems [^1] Weakly-partitive set systems are set systems enjoying some nice closure properties, which were then used to show that some set systems allow canonical tree representations, see for instance the thesis by Montgolfier and Rao [17, 25]. to obtain order-invariant $\operatorname{MSO}$-transductions of modular and split decompositions [12]. It was left as an open question whether one can get rid of the order (see for instance [13] where an overview of the result on hierarchical decompositions was given).
1.1 Our results
@@
In this paper, we consider decompositions given by separations that do not overlap with any other separations of the same type. We view separations of a given kind as a ‘set system’. A set system consists of a set $U$, the universe, and a set $\mathcal{S}$ of subsets of $U$. Two sets overlap if they have non-empty intersection but neither of them is contained in the other. If no two elements in a set system $(U,\mathcal{S})$ overlap, i.e. the set system is laminar, then the subset relation in $(U,\mathcal{S})$ can be described by a rooted tree $T$, called the laminar tree of $(U,\mathcal{S})$. For any set system $(U,\mathcal{S})$ we can look at the subset of strong sets, i.e., sets that do not overlap with any other set, and the laminar tree $T$ they induce. Additionally, we consider systems of bipartitions $\mathcal{B}$ of a universe $U$. Two bipartitions of $U$ overlap if neither side of one of the bipartition is contained in either side of the other bipartition. In case $(U,\mathcal{B})$ has no overlapping bipartitions, then $(U,\mathcal{B})$ can be described by an undirected tree, also called laminar tree, in which bipartitions correspond to edge cuts. We can define the concept of strong bipartitions equivalently and consider the laminar tree induced by the strong bipartitions in $(U,\mathcal{B})$.
@@
Given a graph $G$ we can consider the set system $(V(G),\mathcal{M})$ where $\mathcal{M}$ is the set of all modules in $G$, or the system of bipartition $(V(G),\mathcal{S})$ where $\mathcal{S}$ contains all splits in $G$, or the system of bipartitions $(V(G),\mathcal{B})$ where $\mathcal{B}$ contains all bi-joins in $G$. We obtain the modular decomposition/cotree/split decomposition/bi-join decomposition by equipping the laminar tree of the suitable set system/system of bipartitions with some additional structure. The additional structure allows us to recover the graph from the respective decomposition.
@@
Abstractly, the set systems/systems of bipartitions mentioned are instances of weakly-partitive set systems/weakly-bipartitive systems of bipartitions. Roughly speaking, if a set system $(U,\mathcal{S})$ is well behaved, i.e. $(U,\mathcal{S})$ is weakly-partitive (definition in Section 2), then there is a labelling $\lambda$ of $T$ and a partial order $<$ of its nodes such that $(U,\mathcal{S})$ is completely described by $(T,\lambda,<)$ [7, 25]. A similar statement holds for systems of bipartitions [17].
@@
We show the following, wherein each item $\lambda$ is a suitable labelling of the nodes of the laminar tree $T$, $<$ is a partial ordering of its nodes, and $F$ is an additional edge relation defined only on pairs of siblings in $T$. A visualization how results depend on each other is given in Figure 1.
Theorem 1.1.
There are non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transductions $\tau_{1},\ldots,\tau_{7}$ such that:
  1. 1.
    For any laminar set system $(U,\mathcal{S})$, $\tau_{1}(U,\mathcal{S})$ is non-empty and every output in $\tau_{1}(U,\mathcal{S})$ is a laminar tree $T$ of $(U,\mathcal{S})$ (Theorem 3.1);
  2. 2.
    For any graph $G$, $\tau_{2}(G)$ is non-empty and every output in $\tau_{2}(G)$ is a modular decomposition $(T,F)$ of $G$ (Theorem 4.8);
  3. 3.
    For any cograph $G$, $\tau_{3}(G)$ is non-empty and every output in $\tau_{3}(G)$ is a cotree $(T,\lambda)$ of $G$ (Corollary 4.11);
  4. 4.
    For any graph $G$, $\tau_{4}(G)$ is non-empty and every output in $\tau_{4}(G)$ is a split decomposition $(T,F)$ of $G$ (Theorem 5.11);
  5. 5.
    For any graph $G$, $\tau_{5}(G)$ is non-empty and every output in $\tau_{5}(G)$ is a bi-join decomposition $(T,F)$ of $G$ (Theorem 5.15 and Theorem 5.21);
  6. 6.
@@
    For any weakly-partitive set system $(U,\mathcal{S})$, $\tau_{6}(U,\mathcal{S})$ is non-empty and every output in $\tau_{6}(U,\mathcal{S})$ is a weakly-partitive tree $(T,\lambda,<)$ of $(U,\mathcal{S})$ (Theorem 4.4);
  7. 7.
@@
    For any weakly-bipartitive set system $(U,\mathcal{B})$, $\tau_{7}(U,\mathcal{B})$ is non-empty and every output in $\tau_{7}(U,\mathcal{B})$ is a weakly-bipartitive tree $(T,\lambda,<)$ of $(U,\mathcal{B})$ (Theorem 5.5).
@@
The key step in obtaining these transductions is to transduce the laminar tree $T$ of a set system $(U,\mathcal{S})$. The crux here is to find a suitable representative of each node of $T$ amongst the elements of $U$ and a non-deterministic coloring which allows the assignment of representatives to nodes by means of a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-formula. It should be mentioned that a similar result is claimed in the preprint [3], where a proof sketch designing a $\operatorname{C}_{3}{\!\!}\operatorname{MSO}$-transduction is described. Once the laminar tree is obtained, the additional relations for each decomposition can be obtained using a deterministic $\operatorname{MSO}$-transduction. Notice that for each of these transductions, there exists an inverse deterministic $\operatorname{MSO}$-transduction, namely a transduction that from the tree-like decomposition outputs the original structure.
@@
Section 3Section 4Section 5 Theorem 3.1blablabla laminar set system $\rightarrow$ laminar tree Lemma 4.1blablabla weakly-partitive set system $\rightarrow$ laminar tree Lemma 5.2blablabla weakly-bipartitive set system $\rightarrow$ laminar tree Theorem 4.4blablabla weakly-partitive set system $\rightarrow$ weakly-partitive tree Theorem 4.8blablabla graph $\rightarrow$ modular decomposition Theorem 5.5blablabla weakly-bipartitive set system $\rightarrow$ weakly-bipartitive tree Theorem 5.11blablabla graph $\rightarrow$ split decomposition Corollary 4.11blablabla cograph $\rightarrow$ cotree Theorem 5.15 & 5.21 graph $\rightarrow$ bi-join decomposition
@@
Figure 1: Overview of the various transductions in the paper. An arrow from $x$ to $y$ indicates that result $x$ is used in the proof of result $y$.
1.2 Organization
@@
The paper is organized as follows. In Section 2 we introduce terminology and notation needed. In Section 3 we prove Theorem 3.1. In Section 4 we provide the proof of Theorem 4.4 and Theorem 4.8 and obtain Corollary 4.11. Both formulation of the theorem as well as the proofs of Theorem 5.5, Theorem 5.11, Theorem 5.15 and Theorem 5.21 are postponed to the appendix.
2 Preliminaries
2.1 Graphs, trees, set systems
@@
We use standard terminology of graph theory, and we fix some notations. Given a directed graph $G$, its sets of vertices and edges are denoted by $V(G)$ and $E(G)$, respectively. We denote by uv an edge $(u,v)\in E(G)$. An undirected graph is no more than a directed graph for which $E(G)$ is symmetric (i.e., $uv\in E(G)\iff vu\in E(G)$). The notions of paths, connected components, etc… are defined as usual. Given a subset $Z$ of $V(G)$, we denote by $G[Z]$ the sub-graph of $G$ induced by $Z$.
@@
A tree is a connected undirected graph without cycles. In the context of trees, we use a slightly different terminology than for graphs. In particular, vertices are called nodes, nodes of degree at most $1$ are called leaves, and nodes of degree greater than $1$ are called inner. The set of leaves is denoted $L(T)$; thus the set of inner nodes is $V(T)\setminus L(T)$. For a node $t$ of a tree $T$ and a neighbor $s$ of $t$, we denote by $T^{t}_{s}$ the connected component of $T-t$ containing $s$. We sometimes consider rooted trees, namely trees with a distinguished node, called the root. Rooted trees enjoy a natural orientation of their edges toward the root, which induces the usual notions of parent, child, sibling, ancestor and descendant. Hence, we represent a rooted tree by a set of nodes with an ancestor/descendant relationship (instead of specifying the root). We use the convention that every node is one of its own ancestors and descendants. We refer to ancestors (resp. descendants) of a node that are not the node itself as proper ancestors (resp. proper descendants). For a node $t$ of a rooted tree $T$, we denote by $T_{t}$ the subtree of $T$ rooted in $t$ (i.e., the restriction of $T$ to the set of descendants of $t$).
@@
A set system is a pair $(U,\mathcal{S})$ where $U$ is a finite set, called the universe, and $\mathcal{S}$ is a family of subsets of $U$ where $\emptyset\notin\mathcal{S}$, $U\in\mathcal{S}$, and $\{a\}\in\mathcal{S}$ for every $a\in U$ [^2] Though these restrictions on $\mathcal{S}$ are not usual for set systems, it is convenient for our contribution and it does not significantly impact the generality of set systems: every family $\mathcal{F}$ of subsets of $U$ can be associated with a set system $(U,\mathcal{S})$ where $\mathcal{S}=\big{(}\mathcal{F}\setminus\{\emptyset\}\big{)}\cup\big{\{}U\big{\}}\cup\big{\{}\{a\}\mid a\in U\big{\}}$. Two sets $X$ and $Y$ overlap if they are neither disjoint nor related by containment. A set system $(U,\mathcal{S})$ is said to be laminar (aka overlap-free) when no two sets from $\mathcal{S}$ overlap. By extension, a set family $\mathcal{S}$ is laminar if $\left(\bigcup\mathcal{S},\mathcal{S}\right)$ is a laminar set system (note this also requires that $\emptyset\notin\mathcal{S}$, $\bigcup\mathcal{S}\in\mathcal{S}$, and $\{a\}\in\mathcal{S}$ for every $a\in\bigcup\mathcal{S}$).
A laminar family $\mathcal{S}$ of subsets of $U$ naturally defines a rooted tree where the nodes are the sets from $\mathcal{S}$, the root is $U$, and the ancestor relation corresponds to set inclusion. We call this rooted tree the laminar tree of $U$ induced by $\mathcal{S}$ (or laminar tree of $(U,\mathcal{S})$). In this rooted tree, the leaves are the singletons $\{x\}$ for $x\in U$, which we identify with the elements themselves. That is to say, $L(T)=U$. Laminar trees have the property that each inner node has at least two children. Observe also that the size of a laminar tree is linearly bounded in the size of the universe.
2.2 Logic and transductions

### v4

@@
Abstract.
We give $\operatorname{C}{\!\!}\operatorname{MSO}$-transductions that, given a graph $G$, output its modular decomposition, its split decomposition and its bi-join decomposition. This improves results by Courcelle [Logical Methods in Computer Science, 2006] who gave such transductions using order-invariant $\operatorname{MSO}$, a strictly more expressive logic than $\operatorname{C}{\!\!}\operatorname{MSO}$. Our methods more generally yield $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transductions that output the canonical decompositions of weakly-partitive set systems and weakly-bipartitive systems of bipartitions.
Key words and phrases:
@@ Abstract.
Rutger Campbell: Supported by the Institute for Basic Science (IBS-R029-C1).
Mamadou Moustapha Kanté: Supported by the French National Research Agency (ANR-18-CE40-0025-01 and ANR-20-CE48-0002).
1. Introduction
@@
For a logic $\mathsf{L}$, an $\mathsf{L}$-transduction is a non-deterministic map from a class of relational structures to a new class of relational structures using $\mathsf{L}$-formulas. Transducing a tree-like decomposition is of particular interest. Notably, transducing a decomposition of a graph implies that any property that is definable using a decomposition, is also definable on graphs that admit such a decomposition.To provide an example, a transduction constructing a modular decomposition of a graph $G$ can be used to define a sentence determining that the number of modules in $G$ is even (see Section 4.2 for definitions and Section 4.2 for details). Moreover, tree-like decompositions can be often represented by labelled trees, for which the equivalence of recognisability by a tree automaton and definability in $\operatorname{MSO}$ with modulo counting predicates, denoted $\operatorname{C}{\!\!}\operatorname{MSO}$, is well known [Cou90]. Hence, it is an interesting question to consider what kind of graph decompositions can be transduced using $\mathsf{L}$-transductions for some extension $\mathsf{L}$ of $\operatorname{MSO}$.
@@
In a series of papers [Cou90, Cou91, Cou96, Cou06], Courcelle investigated the relationship between the graph properties that can be defined in an extension of $\operatorname{MSO}$ and the graph properties that can be recognized by a tree automaton where the tree-automaton receives a term representing the input graph. In particular, Courcelle’s theorem states that any graph property that is definable in the logic $\operatorname{C}{\!\!}\operatorname{MSO}$ can be recognized by a tree automaton receiving a tree-decompositions of bounded width [Cou90]. Combining this result with the linear time algorithm for computing tree-decompositions [Bod93], yields that $\operatorname{C}{\!\!}\operatorname{MSO}$ model-checking can be done in linear time on graphs of bounded treewidth. The converse statement – whether recognisability by a tree automaton implies definability in $\operatorname{C}{\!\!}\operatorname{MSO}$ on graphs of bounded treewidth – was conjectured by Courcelle in [Cou90] and finally settled by Bojańczyk and Pilipczuk [BP16]. The key step to obtain this result is obtaining a tree-decomposition of a graph via an $\operatorname{MSO}$-transduction, a strategy which was initially proposed in [Cou91] and is now standard.
@@
The obvious next question is whether an analogous result can be proved for graphs of bounded clique-width and for more general combinatorial objects, most notably, matroids representable over a fixed finite field and of bounded branch-width. Due to the above-mentioned strategy, the key challenge is to produce corresponding tree-like decompositions by $\operatorname{MSO}$-transduction. It is known that clique-width decompositions can be $\operatorname{MSO}$-transduced for graphs of bounded linear clique-width [BGP21]. However, it is unknown if clique-width decompositions can be $\operatorname{MSO}$-transduced in general. In fact, this question remained open even for distance-hereditary graphs, which are precisely graphs of rank-width 1 (thus, of constant clique-width).
@@
Besides tree-decompositions, the problem of transducing cotrees, and in general hierarchical decompositions such as modular decompositions and split decompositions were considered in the literature [Cou96, Cou99, Cou06, Cou13]. In [Cou96], Courcelle provides transductions using order-invariant $\operatorname{MSO}$ for cographs and modular decompositions of graphs of bounded modular width. Order-invariant $\operatorname{MSO}$ allows the use of a linear order of the set of vertices and is more expressive than $\operatorname{C}{\!\!}\operatorname{MSO}$ [GR08]. The applicability of these transductions was later generalized using the framework of weakly-partitive set systems [^1] Weakly-partitive set systems are set systems enjoying some nice closure properties, which were then used to show that some set systems allow canonical tree representations, see for instance the thesis by Montgolfier and Rao [dM03, Rao06]. to obtain order-invariant $\operatorname{MSO}$-transductions of modular and split decompositions [Cou06]. It was left as an open question whether one can get rid of the order (see for instance [Cou12] where an overview of the result on hierarchical decompositions was given).
1.1. Our results
@@
In this paper, we consider decompositions given by nested partitions. We view partitions of a given kind as a ‘set system’. A set system consists of a set $U$, the universe, and a set $\mathcal{S}$ of subsets of $U$. Two sets overlap if they have non-empty intersection but neither of them is contained in the other. If no two elements in a set system $(U,\mathcal{S})$ overlap, i.e. the set system is laminar, then the subset relation in $(U,\mathcal{S})$ can be described by a rooted tree $T$, called the laminar tree of $(U,\mathcal{S})$. For any set system $(U,\mathcal{S})$ we can look at the subset of strong sets, i.e., sets that do not overlap with any other set, and the laminar tree $T$ they induce.
@@
Given a graph $G$ we can consider the set system $(V(G),\mathcal{M})$ where $\mathcal{M}$ is the set of all modules in $G$. We obtain the modular decomposition, or the cotree in case $G$ is a cograph, by equipping the laminar tree of the suitable set system with some additional structure. The additional structure allows us to recover the graph from the respective decomposition.
@@
Abstractly, the set systems mentioned are instances of weakly-partitive set systems. Roughly speaking, if a set system $(U,\mathcal{S})$ is well behaved, i.e. $(U,\mathcal{S})$ is weakly-partitive (definition in Section 2), then there is a tree $T$, a labelling $\lambda$ of $T$ and a partial order << of its nodes such that $(U,\mathcal{S})$ is completely described by $(T,\lambda,<)$ [CHM81, Rao06]. See Figure 1 for an example.
Figure 1. Left: A weakly-partitive set system for which strong sets are indicated by red, thick lines and singletons are omitted. Right: The laminar tree obtain by considering the laminar set system consisting of the strong sets. Here the node labeled $\mathsf{linear}$ plus the linear ordering of its children indicate that the leaves of every << interval corresponds to a set in the set system (e.g. $\{a,b,c\}$ is in the set system while $\{h,f,e,c\}$ is not). Note that some labels of nodes are omitted.
@@
We show the following, wherein each item $\lambda$ is a suitable labelling of the nodes of the laminar tree $T$, << is a partial ordering of its nodes, and $F$ is an additional edge relation defined only on pairs of siblings in $T$. A visualization how results depend on each other is given in Figure 3.
Theorem 1.
There are non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transductions $\tau_{1},\ldots,\tau_{4}$ such that:
  1. (1)
    For any laminar set system $(U,\mathcal{S})$, $\tau_{1}(U,\mathcal{S})$ is non-empty and every output in $\tau_{1}(U,\mathcal{S})$ is a laminar tree $T$ of $(U,\mathcal{S})$ (3);
  2. (2)
@@
    For any weakly-partitive set system $(U,\mathcal{S})$, $\tau_{2}(U,\mathcal{S})$ is non-empty and every output in $\tau_{2}(U,\mathcal{S})$ is a weakly-partitive tree $(T,\lambda,<)$ of $(U,\mathcal{S})$ (15);
  3. (3)
    For any graph $G$, $\tau_{3}(G)$ is non-empty and every output in $\tau_{3}(G)$ is a modular decomposition $(T,F)$ of $G$ (17);
  4. (4)
    For any cograph $G$, $\tau_{4}(G)$ is non-empty and every output in $\tau_{4}(G)$ is a cotree $(T,\lambda)$ of $G$ (19).
Other tree-like graph decompositions can be obtained by considering systems of bipartitions. Aiming to study such decompositions, we can apply our techniques to systems of bipartitions. A systems of bipartitions consists of a universe $U$ and a set $\mathcal{B}$ of bipartitions of $U$. Two bipartitions of $U$ overlap if neither side of one of the bipartition is contained in either side of the other bipartition. In case $(U,\mathcal{B})$ has no overlapping bipartitions, then $(U,\mathcal{B})$ can be described by an undirected tree, also called laminar tree, in which bipartitions correspond to edge cuts. We can define the concept of strong bipartitions equivalently and consider the laminar tree induced by the strong bipartitions in $(U,\mathcal{B})$.
Given a graph $G$, we consider the system of bipartition $(V(G),\mathcal{S})$ where $\mathcal{S}$ contains all splits in $G$, or the system of bipartitions $(V(G),\mathcal{B})$ where $\mathcal{B}$ contains all bi-joins in $G$. Equipping the laminar tree of the respective systems of bipartitions with additional structure yields the split decomposition or bi-join decomposition of the graph. These systems of bipartitions are examples of weakly-bipartitive systems of bipartitions. Similar to weakly-partitive set systems, we can completely describe any weakly-bipartitive system of biparitions by a tree $T$, a labelling $\lambda$ and linear order $_{t}$ of the children of some particular nodes $t$ [dM03]. See Figure 1 for an example.
Figure 2. Left: A weakly-bipartitive system of bipartitions for which strong bipartitions are indicated by red, thick lines and bipartitions of the form $\{\{a\},U\setminus\{a\}\}$ are omitted. Right: The laminar tree obtain by considering the laminar system of bipartitions consisting of the strong bipartitions. The node labeled $\mathsf{linear}$ plus the linear ordering of its children indicate that the leaves of every << interval corresponds to a bipartition in the system of biparitions (e.g. $\{\{a,b,c,i,g\},\{d,e,f,h\}\}$ is in the system of bipartitions while $\{\{a,b,d\},\{c,e,f,g,h,i\}\}$ is not). Note that some labels of nodes are omitted.
We prove the following theorem, where similarly to before each item $\lambda$ is a suitable labelling of the nodes of the laminar tree $T$, << is a partial ordering of its nodes, and $F$ is an additional edge relation defined only on pairs of siblings in $T$.
Theorem 2.
There are non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transductions $\tau_{1},\ldots,\tau_{3}$ such that:
  1. (1)
@@
    For any weakly-bipartitive set system $(U,\mathcal{B})$, $\tau_{1}(U,\mathcal{B})$ is non-empty and every output in $\tau_{1}(U,\mathcal{B})$ is a weakly-bipartitive tree $(T,\lambda,<)$ of $(U,\mathcal{B})$ (22);
  2. (2)
    For any graph $G$, $\tau_{2}(G)$ is non-empty and every output in $\tau_{2}(G)$ is a split decomposition $(T,F)$ of $G$ (25);
  3. (3)
    For any graph $G$, $\tau_{3}(G)$ is non-empty and every output in $\tau_{3}(G)$ is a bi-join decomposition $(T,F)$ of $G$ (26).
@@
The key step in obtaining these transductions is to transduce the laminar tree $T$ of a set system $(U,\mathcal{S})$. The crux here is to find a suitable representative of each node of $T$ amongst the elements of $U$ and a non-deterministic coloring which allows the assignment of representatives to nodes by means of a $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-formula. It should be mentioned that a similar result is claimed in the preprint [Boj23], where a proof sketch designing a $\operatorname{C}_{3}{\!\!}\operatorname{MSO}$-transduction is described. Once the laminar tree is obtained, the additional relations for each decomposition can be obtained using a deterministic $\operatorname{MSO}$-transduction. Notice that for each of these transductions, there exists an inverse deterministic $\operatorname{MSO}$-transduction, namely a transduction that from the tree-like decomposition outputs the original structure. It is worth mentionning that a corollary of 1(2) is a $\operatorname{C}{\!\!}\operatorname{MSO}$-transduction outputting a rank-decomposition of width $1$ for every graph of rank-width $1$.
@@
Section 3Section 4Section 5 3blablabla laminar set system $\rightarrow$ laminar tree 14blablabla weakly-partitive set system $\rightarrow$ laminar tree 21blablabla weakly-bipartitive set system $\rightarrow$ laminar tree 15blablabla weakly-partitive set system $\rightarrow$ weakly-partitive tree 17blablabla graph $\rightarrow$ modular decomposition 22blablabla weakly-bipartitive set system $\rightarrow$ weakly-bipartitive tree 25blablabla graph $\rightarrow$ split decomposition 19blablabla cograph $\rightarrow$ cotree 26 graph $\rightarrow$ bi-join decomposition
@@
Figure 3. Overview of the various transductions in the paper. An arrow from $x$ to $y$ indicates that result $x$ is used in the proof of result $y$.
1.2. Organization
@@
The paper is organized as follows. In Section 2 we introduce terminology and notation needed. In Section 3 we prove 3. In Section 4 we provide the proof of 15 and 17 and obtain 19. In Section 5 we provide the proofs of $2$, 25 and 26.
2. Preliminaries
2.1. Graphs, trees, set systems
@@
We use standard terminology of graph theory, and we fix some notations. We consider graphs to be finite. Given a directed graph $G$, its sets of vertices and edges are denoted by $V(G)$ and $E(G)$, respectively. We denote by uv an edge $(u,v)\in E(G)$. An undirected graph is no more than a directed graph for which $E(G)$ is symmetric (i.e., $uv\in E(G)\iff vu\in E(G)$). The notions of paths, connected components, etc… are defined as usual. Given a subset $Z$ of $V(G)$, we denote by $G[Z]$ the sub-graph of $G$ induced by $Z$.
@@
A tree is a connected undirected graph without cycles. In the context of trees, we use a slightly different terminology than for graphs. In particular, vertices are called nodes, nodes of degree at most $1$ are called leaves, and nodes of degree greater than $1$ are called inner. The set of leaves is denoted $L(T)$; thus the set of inner nodes is $V(T)\setminus L(T)$. For a node $t$ of a tree $T$ and a neighbour $s$ of $t$, we denote by $T^{t}_{s}$ the connected component of TtT-t containing $s$. We sometimes consider rooted trees, namely trees with a distinguished node, called the root. Rooted trees enjoy a natural orientation of their edges toward the root, which induces the usual notions of parent, child, sibling, ancestor and descendant. Hence, we represent a rooted tree by a set of nodes with an ancestor/descendant relationship (instead of specifying the root). We use the convention that every node is one of its own ancestors and descendants. We refer to ancestors (resp. descendants) of a node that are not the node itself as proper ancestors (resp. proper descendants). For a node $t$ of a rooted tree $T$, we denote by $T_{t}$ the subtree of $T$ rooted in $t$ (i.e., the restriction of $T$ to the set of descendants of $t$).
@@
A set system is a pair $(U,\mathcal{S})$ where $U$ is a finite set, called the universe, and $\mathcal{S}$ is a family of subsets of $U$ where $\emptyset\notin\mathcal{S}$, $U\in\mathcal{S}$, and $\{a\}\in\mathcal{S}$ for every $a\in U$. [^2] Though these restrictions on $\mathcal{S}$ are not usual for set systems, it is convenient for our contribution and it does not significantly impact the generality of set systems: every family $\mathcal{F}$ of subsets of $U$ can be associated with a set system $(U,\mathcal{S})$ where $\mathcal{S}=\big{(}\mathcal{F}\setminus\{\emptyset\}\big{)}\cup\big{\{}U\big{\}}\cup\big{\{}\{a\}\mid a\in U\big{\}}$. Two sets $X$ and $Y$ overlap if they are neither disjoint nor related by containment. A set system $(U,\mathcal{S})$ is said to be laminar (aka overlap-free) when no two sets from $\mathcal{S}$ overlap. By extension, a set family $\mathcal{S}$ is laminar if $\left(\bigcup\mathcal{S},\mathcal{S}\right)$ is a laminar set system (note this also requires that $\emptyset\notin\mathcal{S}$, $\bigcup\mathcal{S}\in\mathcal{S}$, and $\{a\}\in\mathcal{S}$ for every $a\in\bigcup\mathcal{S}$).
A laminar family $\mathcal{S}$ of subsets of $U$ naturally defines a rooted tree where the nodes are the sets from $\mathcal{S}$, the root is $U$, and the ancestor relation corresponds to set inclusion, i.e., nodes corresponding to sets $S\subset S^{\prime}\in\mathcal{S}$ are adjacent in the tree if there is no proper superset of $S$ in $\mathcal{S}$ that is a proper subset of $S^{\prime}$. We call this rooted tree the laminar tree of $U$ induced by $\mathcal{S}$ (or laminar tree of $(U,\mathcal{S})$). In this rooted tree, the leaves are the singletons $\{x\}$ for $x\in U$, which we identify with the elements themselves. That is to say, $L(T)=U$. Laminar trees have the property that each inner node has at least two children due to the definition of set systems demanding each singleton being in the set system. Observe also that the size of a laminar tree is linearly bounded in the size of the universe.
2.2. Logic and transductions